a stone is dropped from the roof of building takes 5 sec to reach the ground.The sound of striking is heard after 5.4 sec after the stone is dropped.The velocity of sound is?
 

Grade:11

4 Answers

Aniket Tiwari
6 years ago
v= ut + at sq
v= 1/2* 9.8*5*5
s= v sq/2g
i think also that the question is is also incomplite u should check the question again because u didnt had mention the hight of the roof from ground than also i had answer some parts from it sq means square of the given value
swarup kumar majee
6 years ago
u=0,t=5sec,a=9.8m/s2
so ,hieght of the building is h=1/2at2
h=122.5m
sound is heard after 0.4sec after the sound is produced 
so velocity of sound is=122.5/0.4=306.25 m/s
 
 
Jefina Mathew
6 years ago
Here u=0,t=5s,a=10m/s^2S=1/2gt^2S=125mSound is heard after 0.4sVelocity of sound = distance travelled/timeVelocity=125/0.4=312.5
Nasrin
6 years ago
We know s=ut+1/2at^2.putting u=0,a=10 and t=5 we get s=125m.so the building is 125m high.Sound is heard after 5.4-5=0.4 s.Again we know, velocity *time=distanceSo, v*0.4=125,v=125/0.4=312.5